OpenStudy (mortonsalt):

I have a question on Lagrange Multipliers: if f(x,y)=e^(-xy) and the constraint is x^2+4y^2-1<=0, how do I find the critical points when the case is x^2+4y^2-1<0?

OpenStudy (mortonsalt):

i.e. \[x^2+4y^2-1<0\] I've already found the critical points for when \[x^2+4y^2-1=0\] And I know that df/dx=-ye^(-xy) and df/dy=(-xe^(-xy))

OpenStudy (irishboy123):

do your normal equality constraint stuff on the ellipse, which seems you have done and then look for a CP, ie \(f_x = f_y = 0\), within the ellipse. you could luck out if, say, there was a local max or a min in there.... then you know that's what you are looking for, optimisation-wise. but here, **think** you'll find a saddle point at (0,0), eg note \(f(x,y) = 1\) all along the x and y axes as well as at origin......and you have some kind of symmetry going on as the exponent is \(-xy\) which will be -ve in Q1, Q3 and vice versa in others i'm getting \(e^{1/4}\) (1.28) and \(e^{-1/4}\) (0.78) on the ellipse itself ie along \(y = \pm {1 \over 2} x\)

OpenStudy (mathmate):

@IrishBoy123 I get four solutions that satisfy the Lagrange conditions, \(x=\pm1/\sqrt 2, y=\pm x/2,\lambda=e^{1/4}/4\) and \(x=\pm1/\sqrt 2, y=\mp x/2,\lambda=-e^{-1/4}/4\) all of which are for equality (=0) in the constraint. I'm not sure how to take care of the <0 case(s). Any ideas?

OpenStudy (irishboy123):

@mathmate first of all, there is a way to solve Lagrange Multiple type optimisations with inequalities in the constraints. It involves the use of **slack variables**. i find it absolutely horrible. It seems to me to undo the point of the method, which is simplified algebra, by adding more variables. i don't know if you know that approach, i have never finished one off. so inside the constraint, i see no problem in just looking for fixed points. if you find a max or min inside the ellipse, then that is a local max or min for the ellipse and its region. job done. here, there is no max or min. \(f_x = f_y = 0\) occurs at the origin. but it is a saddle. so its not a max or a min so forget about it! i can't remember if i looked at the Hessian when i first did this but what you can do here, because the function is so clean, is you can draw contour lines. so the line \(y = 1/x\) has \(f(x,y) = 1/e\), the line \(y = 2/x\) has \(f(x,y) = 1/e^2\) etc . so \(f(0,0) = 1\) and the function grows exponentially in Q2 & Q4 and dies exponentially in Q1 & Q3 . sure you know all that, just saying. you can then see how the saddle is symmetric and in particular, because it is not a max or min, and so can be ignored. not sure if that helps. btw, i agree with location of your fixed points on the ellipse but i don't get the 4 denominator when i plug those back into \( f(x,y) \) to get actual values.

OpenStudy (mathmate):

@irishboy123 Yes, it is a saddle point. I had it plotted and it's quite visual. http://prnt.sc/c2pxb2 Even though (0,0) satisfies constraints, it would not be the point we're looking for. The four points I gave when substituted back into f(x,y) all gave e^(-1/4) (approx. 0.7788), which are less that f(0,0)=1. As a check, I solved for x^2+4y^2=1/2 and =1/4, both of which gave values between 0.7788 and 1, probably evident from the fact that the constraint is an ellipse, so f(x,y) along y=\(\pm\)x/2 is expected to be monotonic. It's unfortunate that Lagrange multipliers handles inequality constraints with difficulty. Yes, I understand the role of slack variables from linear programming. It might be worth a try for me, some time down the road! Thank you for the detailed explanations.

OpenStudy (irishboy123):

brilliant! thank you.

OpenStudy (thewafflebro):

wwow

OpenStudy (irishboy123):

@mathmate i just remembered something else as well - The Extreme Value Theorem is the proper name given to this idea, of just looking for CP's in the region I'm sure you know that already. but i thought i'd mention.

OpenStudy (mathmate):

@irishboy123 Thank you, that's a good reminder. Sometimes we do things by reflex. lol

OpenStudy (leenathan):

...

OpenStudy (mortonsalt):

@IrishBoy123 @mathmate Hey guys! I'm so sorry I just got back to this now. Thank you so much, that helps out a lot.

OpenStudy (jsnake789):

first of all, there is a way to solve Lagrange Multiple type optimisations with inequalities in the constraints. It involves the use of **slack variables**. i find it absolutely horrible. It seems to me to undo the point of the method, which is simplified algebra, by adding more variables. i don't know if you know that approach, i have never finished one off. so inside the constraint, i see no problem in just looking for fixed points. if you find a max or min inside the ellipse, then that is a local max or min for the ellipse and its region. job done. here, there is no max or min. fx=fy=0 occurs at the origin. but it is a saddle. so its not a max or a min so forget about it! i can't remember if i looked at the Hessian when i first did this but what you can do here, because the function is so clean, is you can draw contour lines. so the line y=1/x has f(x,y)=1/e, the line y=2/x has f(x,y)=1/e2 etc . so f(0,0)=1 and the function grows exponentially in Q2 & Q4 and dies exponentially in Q1 & Q3 . sure you know all that, just saying. you can then see how the saddle is symmetric and in particular, because it is not a max or min, and so can be ignored. not sure if that helps. btw, i agree with location of your fixed points on the ellipse but i don't get the 4 denominator when i plug those back into f(x,y) to get actual values.

OpenStudy (mathmate):

@JSnake789 Yes, your comments pretty much confirm what @irishboy123 expressed earlier that the use of slack variables is counter-nature for inequality constraints. Well, all four solutions give f(x,y)=\(e^{-1/4}\) or 0.7788. The value of \(\lambda\) is not used in calculating f(x,y), just for the optimization part. Or perhaps I misunderstood your comment about the denominators.

umerlodhi (umerlodhi):

@satellite73 can you warn me? plzzz

OpenStudy (leenathan):

:3

OpenStudy (vanesaretana):

:3

OpenStudy (ravegex):

.-.

OpenStudy (thewafflebro):

────────▌▒█───────────▄▀▒▌ ────────▌▒▒▀▄───────▄▀▒▒▒▐ ───────▐▄▀▒▒▀▀▀▀▄▄▄▀▒▒▒▒▒▐ ─────▄▄▀▒▒▒▒▒▒▒▒▒▒▒█▒▒▄█▒▐ ───▄▀▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▀██▀▒▌ ──▐▒▒▒▄▄▄▒▒▒▒▒▒▒▒▒▒▒▒▒▀▄▒▒▌ ──▌▒▒▐▄█▀▒▒▒▒▄▀█▄▒▒▒▒▒▒▒█▒▐ ─▐▒▒▒▒▒▒▒▒▒▒▒▌██▀▒▒▒▒▒▒▒▒▀▄▌ its raining Doges... ─▌▒▀▄██▄▒▒▒▒▒▒▒▒▒▒▒░░░░▒▒▒▒▌ ─▌▀▐▄█▄█▌▄▒▀▒▒▒▒▒▒░░░░░░▒▒▒▐ ▐▒▀▐▀▐▀▒▒▄▄▒▄▒▒▒▒▒░░░░░░▒ ▐▒▒▒▀▀▄▄▒▒▒▄▒▒▒▒▒▒░░░░░░▒▒▒▐ ─▌▒▒▒▒▒▒▀▀▀▒▒▒▒▒▒▒▒░░░░▒▒▒▒▌ ─▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▐ ──▀▄▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▄▒▒▒▒▌ ────▀▄▒▒▒▒▒▒▒▒▒▒▄▄▄▀▒▒▒▒▄▀ ───▐▀▒▀▄▄▄▄▄▄▀▀▀▒▒▒▒▒▄▄▀ ──▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▀

OpenStudy (redheadangel):

I have a really easy question!!

OpenStudy (redheadangel):

Would one of y'all help me? Given the equation, y = 3(0.87)^x , where y is the amount of money in thousands in x years, describe what is happening? The amount of money is decreasing at a rate of 13% per year The amount of money is decreasing at a rate of 87% per year The amount of money is decreasing at a rate of 13 per year The amount of money is decreasing at a rate of 87 per year

OpenStudy (mathmate):

@redheadangel Please post a new question, since this question does not relate to the original post.

OpenStudy (dallascowboys88):

...??